My problem is this 4(x+2)+2=4x+2(-2-x)+4
4x+8+2=4x-4-2x+4
4x+10=2x
2x=10
x=5
But when I check my answer, they don't equal. What I'm I doing wrong?

Answer by checkley71(8403)

4(X+2)+2=4X+2(-2-X)+4
4X+8+2=4X-4-2X+4
4X-4X+2X=-8-2-4+4
2X=-10 YOU DIDN'T CHANGE THE SIGN OF 10 WHEN YOU TRANSFERED IT TO THE RIGHT SIDE OF THE EQUATION.
X=-10/2
X=-5